OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

Solve for x by dividing both sides of the equation by -4. |x + 5| = 4 with absolute value there may be more than one answer so set up the problem like this -4 = |x + 5| = 4 Now you have two problems. Subtract 5 from both sides -4 - 5 = x x = 4 - 5 -9 = x and x = -1

OpenStudy (anonymous):

Thank you!